Two resistors are connected in parallel across an ideal 12.0 V battery. Resistor A has a value of 28 Ω and resistor B carries a current of 0.2505 A. What is the total power dissipated in the circuit?Two resistors are connected in parallel across an ideal 12.0 V battery. Resistor A has a value of 28 Ω and resistor B carries a current of 0.2505 A. What is the total power dissipated in the circuit?
Answer #1

power in resistor A = V2/R = 144/28 = 5.143 W

power in resistor B = VI = 12 x 0.2505 = 3.006 W

total power = 5.143 + 3.006 = 8.149 W